# Quantum Optimization Benchmark Library The Intractable Decathlon

### v1

@@ 4.8.2 Problem statement
We can formulate this network design problem as a multicommodity design problem as follows: For $k,i,j\in\{1,\ldots,n\}$ we define binary variables $x_{ij}\in\{0,1\}$ which are $1$ iff an arc is established between vertex $i$ and $j$ and 00, otherwise. Further, we define non-negative integer variables $f_{kij}\in\mathbb{Z}_{+}$ that represent the flow of commodity $k$ between vertices $i$ and $j$, and a suitable big constant $M$. Then we get:
@@
subject $\text{subject to}\sum\limits_{j\neq i}x_{ij}$ $=p$ for all $\text{for all }i\>,$ (13)
$\sum\limits_{j\neq i}x_{ji}$ $=p$ for all $\text{for all }i\>,$ (14)
$f_{kij}$ $\leq Mx_{ij}$ for all $\text{for all }k,i\neq j\>,$ (15)
$\sum\limits_{j\neq i}f_{kij}-\sum\limits_{j\neq i}f_{kji}$ $=t_{ki}$ for all $\text{for all }i,k\>,$ (16)
$\sum\limits_{k}f_{kij}$ $\leq z$ for all $\text{for all }i\neq j\>.$ (17)
@@ 4.8.4 Classical Baseline
We provide a basic IP formulation of the problem. More sophisticated cutting planes are described in [21]. Results with the basic IP formulation found using Gurobi 11.0.0 on an AMD EPYC 7542 32-core processor with 64 threads and a time limit of 2 hours are summarized in Figure 14. We were able to solve the problem up to size $n=9$ to proven optimality within the time limit. For larger sizes, heuristic, but not necessarily optimal, solutions are provided. The instance with 24 vertices and continuous flow variables is also part of MIPLIB2017 [84], the latest results of which can be found at [53]. Note that those results correspond to the unscaled model version with continuous variables, which, therefore, has a slightly different objective value than the scaled one. However, the solutions are compatible.
Figure 14: (Network) Runtime is the time to compute a proven optional solution. Time-to-solution is the time until the best solution was found. Rel. opt. Gap is the gap to the optimal or best-known solution. MIP Gap is the gap to the lower bound. Only gaps $>0$ are plotted.
@@ 4.9.1 Background
The Vehicle Routing Problem (VRP) is the problem of determining the most efficient routes for a fleet of vehicles to deliver goods or services to a set of customers, typically starting and ending at a central depot. This problem is highly relevant for many real-world applications in logistics, transportation, and supply chain management, see, e.g., [41]. Originally proposed in [61] as ‘the truck dispatching problem’ many different variants of VRP have emerged since then [101, 234]. Providing a collection of relevant and difficult VRP instances is a complex task, because there are numerous variants. Although classical heuristics for VRP are practically very successful, many papers on quantum approaches for VRP problems have been published [246, 160, 144, 72]. We include the VRP problem class in this collection to facilitate a comparison with existing work and to explore the capabilities of quantum optimization algorithms for a practically relevant problem.
We limit ourselves to a specific variant of VRP, the Capacitated Vehicle Routing Problem (CVRP), where the objective is to minimize the aggregated costs across all routes while satisfying the vehicle capacity, effectively limiting the number of goods that each vehicle can carry. This variant, in fact, corresponds to the original problem formulation from [61], where it was presented as a generalization of the TSP [177]. Compared to TSP, VRP has an additional level of decision-making, as customers not only have to be served but also allocated to vehicles. An example CVRP instance and its optimal solution are visualized in Figure 15.
Why is it interesting? Determining the optimal solution to a Vehicle Routing Problem (VRP) is NP-hard [148] and of high practical relevance as it has many direct industrial applications and, hence, the potential for a significant business value impact. Similar to the network problem from the previous section, it is trivial to find feasible (but not necessarily optimal) solutions. Due to the complexity of the problem, mostly small instances can be solved exactly, but heuristics and metaheuristics are often well suited to tackle real-world instances. This problem class has already received significant attention from the quantum community.
@@ 4.9.2 Problem statement
Consider a CVRP with $n\geq 1$ customers, where each customer $i$ has a demand $d_{i}\in\mathbb{N}$ with $i\in\{1,\dots,n\}$. Furthermore, we assume that there is a homogeneous fleet of $K\geq 1$ vehicles, where each vehicle has a total capacity of $Q\in\mathbb{N}$. Each vehicle starts and ends its tour at the depot. The cost of moving a vehicle from customer $i$ to $j$ is given by $c_{ij}\in\mathbb{N}$ with $i,j\in\{0,\dots,n\}$ and $i\neq j$, where “customer 00” is used to represent the depot. With this definition, we presume all-to-all connectivity between the customers (and the depot). While the cost matrix is generally arbitrary, we focus here on a special case relevant to the benchmarking instances we provide, where the costs are based on graph vertex distances in a metric space that satisfies non-negativity (as given by the cost domain) and symmetry (i.e., $c_{ij}=c_{ji}\,\forall\,i,j\in\{0,\dots,n\}$, the condition for the aforementioned symmetric CVRP), which means they also obey the triangle inequality (i.e., $c_{ij}\leq c_{ik}+c_{kj}\,\forall\,i,j,k\in\{0,\dots,n\}$). Each customer must be served exactly once by a vehicle, while the total customer demand on a given route must not exceed the capacity limit of the vehicle. The optimization goal is to determine a feasible solution that minimizes the aggregated costs across all routes. This does not necessarily require the use of all $K$ vehicles, i.e., some may remain in the depot and, hence, do not contribute to the costs.
@@
subject $\text{subject to}\sum_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{n+1}x_{ij}$ $=1$ for all $\text{for all }i\in\{1\ldots,n\}\>,$ (20)
$\sum_{\begin{subarray}{c}i=0\\ i\neq h\end{subarray}}^{n}x_{ih}$ $=\sum_{\begin{subarray}{c}j=1\\ j\neq h\end{subarray}}^{n+1}x_{hj}$ for all $\text{for all }h\in\{1,\ldots,n\}\>,$ (21)
@@
$y_{j}$ $\geq y_{i}+d_{j}x_{ij}-Q(1-x_{ij})$ for all $\text{for all }i\neq j\in\{0,\ldots,n+1\}\>,$ (23)
$d_{i}$ $\leq y_{i}\leq Q,$ for all $\text{for all }i\in\{0,\ldots,n+1\}\>,$ (24)
$x_{ij}$$\in\{0,1\}$ for all $\text{for all }i,j\in\{0,\ldots,n+1\}\>,$ (25)
$y_{i}$ $\in\mathbb{N}_{0}$ for all $\text{for all }i\in\{0,\ldots,n+1\}\>.$ (26)
In this IP, the optimization is done with respect to two types of decision variables. First, the binary variables $x\in\{0,1\}^{(n+2)\times(n+2)}$ and second, the integer variables $y\in\mathbb{N}_{0}^{n+2}$. The binary variable $x_{ij}$ takes the value 1 if and only if there is a direct route between customers $i$ and $j$ for $i,j\in\{0,\dots,n+1\}$, where “customer 00” is the depot, as introduced above. To simplify the notation, we also symbolically introduce “customer $n+1$” to represent the same depot, which allows us to distinguish between starting at the depot ($i=0$) and ending at the depot ($i=n+1$), where $c_{j,n+1}:=c_{0j}$ for all $j\in\{1,\dots,n\}$. Furthermore, the integer variable $y_{i}$ represents the cumulated demand on the route that visits customer $i$ for $i\in\{0,\dots,n+1\}$, where $d_{0}:=d_{n+1}:=0$.
@@ 4.9.3 Instances
A survey on benchmark data sets can be found in [86]. All instances in the repository are given for $k=4$ vehicles and $n=20$ customers and were created to be 100% tight, i.e., the sum of the demands equals the vehicle capacity $Q$. The instances follow the TSPLIB/CVRPLIB format [151]. We randomly generated and solved several thousand instances and selected those that took the longest to solve, which is still only seconds on modern digital systems. Out of the set of problem sizes considered, the chosen instances are among the most difficult. It has proven unexpectedly hard to find small-to-medium instances that are primal difficult. Classical methods solve these sizes quickly, see [223, 86]. This indicates that quantum approaches would have to be extremely fast to be competitive on these problems in the near-term. We plan to add and investigate larger instances in the future.
@@ 4.9.4 Classical Baseline
Table 8 shows the optimal cost and the average time taken by HGS-CVRP [102, 232] to find this solution, measured over 10 runs with different seeds and a time limit of 10 seconds. Notably, the optimal solution could not be found within this limit, for instance, 21. However, given that HGS-CVRP takes only a few milliseconds to find an optimal solution for typical instances with $n=20$, we expect that it would be found with a marginally longer time limit. The optimal solutions were certified using VRPSolver [70, 189, 237, 238].
@@ 4.10.1 Background
Notably, the name of the problem is given by the two fixed parameters. The ODiP has received little attention to date. The DDP has been considered extensively, with the foundations discussed already in the literature from the 1960s [110, 39], and further explored in the 1980s, e.g., [43, 44]. For a comprehensive survey, we refer the reader to [176]. In this work, we consider the ODP, for which–to our knowledge–an NP-hardness proof is not available in the open literature. The ODP arises in the search for network topologies with low latency [216], and a variety of methodologies to construct networks with low latency are discussed in the literature [23, 208]. We use the ODP variant used in the GraphGolf competitions [127]: the goal is to solve the ODP, using the average shortest path length (ASPL) as a tie-breaker for graphs with the same diameter, with lower ASPL being preferred. This problem is difficult to solve in practice [127].
@@ 4.10.2 Problem statement
Be $\Delta(s,t)$, $s,t\in V$ the shortest distance of two vertices in a simple unweighted graph $G=(V,E)$. Given a number $n\in\mathbb{N}$ of vertices, and a maximum vertex degree $d\in\mathbb{N}$. We are looking for a simple graph $G=(V,E)$, $V=\{1,\ldots,n\}$, $E\subseteq\{(u,v)\in V\times V|u\neq v\}$ with $\delta(v)\leq d$, for all $v\in V$, that minimizes $\max_{s,t\in V}\Delta(s,t)$. Multiple formulations for the ODP are possible; for an overview, we refer to [239]. There is considerable freedom in the choice of how the lengths of the all-pairs shortest paths are computed.
@@ 4.10.3 Instances
The instances range from $15$ vertices and degree $3$ to $10^{6}$ vertices and degree $32$. A detailed presentation of the provided instances can be found in Table 9. We provided a program to check solutions, i.e., given a triple $(n,d,k)$ and a graph, it verifies that the graph has $n$ nodes, a maximum degree of $d$, and a diameter less or equal to $k$.
Figure 16: (Topology) Absolute MIP gap corresponding to order and degree of the flow (top) / quadratic (bottom left) and linearized (bottom right) Seidel model. The size of the points indicates the running time–bigger is longer. Furthermore, the timeout is set to 7200s.
@@ 4.10.4 Classical Baseline
Our baseline was computed using Gurobi 11.0.0 on an AMD EPYC 7542 32-core processor using 64 threads with a time limit of 7200s. In the repository, we provide three different MIP formulations. The first model uses a standard s-t flow model to compute shortest path values. The second model uses Seidel’s quadratic all-pairs shortest path formulation. The third model is a linearization of Seidel’s quadratic model. This problem is hard to model as a compact integer program. Gaps and runtimes can be found in Figure 16. The results indicate that we can prove optimality only on rather small instances if we use an MIP solver. With the provided models, even solving $n=30,d=6$ takes too long. However, as can be seen on [127], using heuristic methods, it is possible to compute solutions to larger instances.
@@ 5 Submission Guidelines and Illustrative Results
The following section presents guidelines for submitting benchmark results to QOBLIB and provides exemplary baseline results for three selected problem classes–Low Autocorrelation Binary Sequences, Minimum Birkhoff Decomposition and Independent Set problem–each solved with another quantum optimization algorithm. These results can also be found in the QOBLIB repository [193].
We provide a submission template in Table 10. The metrics were chosen to enable a fair and systematic comparison of solutions and algorithms. More specifically, we ask for the identification of the concrete problem instance and the submitter(s). Given that not every detail can be given in a compact form, submissions ideally refer to a paper or a code repository with further information, such as hyperparameters, additional hardware specifications, software versions, etc. Furthermore, the best objective value found (for optimization problems) as well as the corresponding solution (in addition to the template) should be given. If supported by an algorithm, an a posteriori bound on the optimal objective value can be provided. Next, the chosen modeling approach should be described (e.g., QUBO, ILP, etc.) as well as the resulting number and type of decision variables and (non-zero) coefficients needed to represent the considered problem instance. Further, the submission should briefly summarize the complete optimization workflow to facilitate reproducability. This includes pre-processing, pre-solver, main optimization algorithm, and post-processing, as well as an indication if the algorithm is deterministic or stochastic. Stochastic algorithms are generally recommended to be repeated multiple times. In this case, the number of successful runs that result in feasible solutions or return solutions close to the best found solutions (cf. Table 10 for more details) should be reported. Last, the overall runtime and the runtime spent on the various hardware platforms should be provided. In case of multiple repetitions, the average runtime over all repetitions may be reported. Additional information, such as the distribution of the runtimes or potential correlations with solution quality, is encouraged to be described in a corresponding publication or reference. As discussed in Section 3, the runtimes should not include potential queuing times for hardware access. For each problem, we provide solution checkers in [193]. These should be used to verify proposed solutions before a benchmark is submitted to QOBLIB.
@@
Decision Variables Range Range of the decision variables, i.e., min/max values.
@@
Coefficients Type Type of coefficients such as integer, binary, continuous.
@@
Workflow Description of the optimization workflow: pre-processing, pre- solvers, optimization algorithms, and post-processing, etc.
@@
Total Runtime Total runtime to run the complete workflow.
@@
Table 10: The table summarizes the metrics that should be reported when submitting benchmark results to QOBLIB. All runtimes should be reported as average if multiple algorithm runs were executed.

### v2

@@ 4.8.2 Problem statement
We can formulate this network design problem as a multicommodity design problem as follows: For $k,i,j\in\{1,\ldots,n\}$ we define binary variables $x_{ij}\in\{0,1\}$ which are $1$ iff an arc is established between vertex $i$ and $j$ and 0, otherwise. Further, we define non-negative integer variables $f_{kij}\in\mathbb{Z}_{+}$ that represent the flow of commodity $k$ between vertices $i$ and $j$, and a suitable big constant $M$. Then we get:
@@
$\text{subject to}\sum\limits_{j\neq i}x_{ij}$ $=p$ $\text{for all }i\>,$ (13)
$\sum\limits_{j\neq i}x_{ji}$ $=p$ $\text{for all }i\>,$ (14)
$f_{kij}$ $\leq Mx_{ij}$ $\text{for all }k,i\neq j\>,$ (15)
$\sum\limits_{j\neq i}f_{kij}-\sum\limits_{j\neq i}f_{kji}$ $=t_{ki}$ $\text{for all }i,k\>,$ (16)
$\sum\limits_{k}f_{kij}$ $\leq z$ $\text{for all }i\neq j\>.$ (17)
@@ 4.8.4 Classical Baseline
We provide a basic IP formulation of the problem. More sophisticated cutting planes are described in [22]. Results with the basic IP formulation found using Gurobi 11.0.0 on an AMD EPYC-7542 32-core processor with 64 threads and a time limit of 2 hours are summarized in Figure 14. We were able to solve the problem up to size $n=9$ to proven optimality within the time limit. For larger sizes, heuristic, but not necessarily optimal, solutions are provided. The instance with 24 vertices and continuous flow variables is also part of MIPLIB2017 [89], the latest results of which can be found at [58]. Note that those results correspond to the unscaled model version with continuous variables, which, therefore, has a slightly different objective value than the scaled one. However, the solutions are compatible.
Figure 14: (Network) Runtime is the time to compute a proven optimal solution. Time-to-solution is the time until the best solution was found. Rel. opt. Gap is the gap to the optimal or best-known solution. MIP Gap is the gap to the lower bound. Only gaps >0>0 are plotted.
@@ 4.9.1 Background
The Vehicle Routing Problem (VRP) is the problem of determining the most efficient routes for a fleet of vehicles to deliver goods or services to a set of customers, typically starting and ending at a central depot. This problem is highly relevant for many real-world applications in logistics, transportation, and supply chain management, see, e.g., [45]. Originally proposed in [66] as ‘the truck dispatching problem’, many different variants of VRP have emerged since then [108, 243]. Providing a collection of relevant and difficult VRP instances is a complex task because there are numerous variants. Although classical heuristics for VRP are practically very successful, many papers on quantum approaches for VRP problems have been published [256, 169, 152, 77]. We include the VRP problem class in this collection to facilitate a comparison with existing work and to explore the capabilities of quantum optimization algorithms for a practically relevant problem.
We limit ourselves to a specific variant of VRP, the Capacitated Vehicle Routing Problem (CVRP), where the objective is to minimize the aggregated costs across all routes while satisfying the vehicle capacity, effectively limiting the number of goods that each vehicle can carry. This variant, in fact, corresponds to the original problem formulation from [66], where it was presented as a generalization of the TSP [186]. Compared to TSP, VRP involves an additional level of decision-making, as customers not only need to be served but also assigned to vehicles. An example CVRP instance and its optimal solution are visualized in Figure 15.
Why is it interesting? Determining the optimal solution to a Vehicle Routing Problem (VRP) is NP-hard [156] and of high practical relevance as it has many direct industrial applications and, hence, the potential for a significant business value impact. Similar to the network problem from the previous section, it is trivial to find feasible (but not necessarily optimal) solutions. Due to the complexity of the problem, mostly small instances can be solved exactly, but heuristics and metaheuristics are often well-suited to tackle real-world instances. This problem class has already received significant attention from the quantum community.
@@ 4.9.2 Problem statement
Consider a CVRP with $n\geq 1$ customers, where each customer $i$ has a demand $d_{i}\in\mathbb{N}$ with $i\in\{1,\dots,n\}$. Furthermore, we assume that there is a homogeneous fleet of $K\geq 1$ vehicles, where each vehicle has a total capacity of $Q\in\mathbb{N}$. Each vehicle starts and ends its tour at the depot. The cost of moving a vehicle from customer $i$ to $j$ is given by $c_{ij}\in\mathbb{N}$ with $i,j\in\{0,\dots,n\}$ and $i\neq j$, where “customer 0” is used to represent the depot. With this definition, we presume all-to-all connectivity between the customers (and the depot). While the cost matrix is generally arbitrary, we focus here on a special case relevant to the benchmarking instances we provide, where the costs are based on graph vertex distances in a metric space that satisfies non-negativity (as given by the cost domain) and symmetry (i.e., $c_{ij}=c_{ji}\,\forall\,i,j\in\{0,\dots,n\}$, the condition for the aforementioned symmetric CVRP), which means they also obey the triangle inequality (i.e., $c_{ij}\leq c_{ik}+c_{kj}\,\forall\,i,j,k\in\{0,\dots,n\}$). Each customer must be served exactly once by a vehicle, while the total customer demand on a given route must not exceed the capacity limit of the vehicle. The optimization goal is to determine a feasible solution that minimizes the aggregated costs across all routes. This does not necessarily require the use of all $K$ vehicles, i.e., some may remain in the depot and, hence, do not contribute to the costs.
@@
$\text{subject to}\sum_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{n+1}x_{ij}$ $=1$ $\text{for all }i\in\{1\ldots,n\}\>,$ (20)
$\sum_{\begin{subarray}{c}i=0\\ i\neq h\end{subarray}}^{n}x_{ih}$ $=\sum_{\begin{subarray}{c}j=1\\ j\neq h\end{subarray}}^{n+1}x_{hj}$ $\text{for all }h\in\{1,\ldots,n\}\>,$ (21)
@@
$y_{j}$ $\geq y_{i}+d_{j}x_{ij}-Q(1-x_{ij})$ $\text{for all }i\neq j\in\{0,\ldots,n+1\}\>,$ (23)
$d_{i}$ $\leq y_{i}\leq Q,$ $\text{for all }i\in\{0,\ldots,n+1\}\>,$ (24)
$x_{ij}$ $\in\{0,1\}$ $\text{for all }i,j\in\{0,\ldots,n+1\}\>,$ (25)
$y_{i}$ $\in\mathbb{N}_{0}$ $\text{for all }i\in\{0,\ldots,n+1\}\>.$ (26)
In this IP, the optimization is done with respect to two types of decision variables. First, the binary variables $x\in\{0,1\}^{(n+2)\times(n+2)}$ and second, the integer variables $y\in\mathbb{N}_{0}^{n+2}$. The binary variable $x_{ij}$ takes the value 1 if and only if there is a direct route between customers $i$ and $j$ for $i,j\in\{0,\dots,n+1\}$, where “customer 0” is the depot, as introduced above. To simplify the notation, we also symbolically introduce “customer n+1n+1” to represent the same depot, which allows us to distinguish between starting at the depot ($i=0$) and ending at the depot ($i=n+1$), where $c_{j,n+1}:=c_{0j}$ for all $j\in\{1,\dots,n\}$. Furthermore, the integer variable $y_{i}$ represents the cumulated demand on the route that visits customer $i$ for $i\in\{0,\dots,n+1\}$, where $d_{0}:=d_{n+1}:=0$.
@@ 4.9.3 Instances
A survey on benchmark data sets can be found in [92]. All instances in the repository are given for $k=4$ vehicles and $n=20$ customers and were created to be 100% tight, i.e., the sum of the demands equals the vehicle capacity $Q$. The instances follow the TSPLIB/CVRPLIB format [161]. We randomly generated and solved several thousand instances, selecting those that took the longest to solve, which is still only seconds on modern digital systems. Out of the set of problem sizes considered, the chosen instances are among the most difficult. It has proven unexpectedly hard to find small-to-medium instances that are primal difficult. Classical methods solve these sizes quickly, see [232, 92]. This indicates that quantum approaches would have to be extremely fast to be competitive on these problems in the near term. We plan to add and investigate larger instances in the future.
@@ 4.9.4 Classical Baseline
It should be noted that numerous open-source codes for VRP are available, e.g., [157, 99, 109], with some even utilizing GPU acceleration [55]. Additionally, several companies offer specialized, highly optimized solvers for this problem; however, it is difficult to assess their performance. Furthermore, they are tailored to real-world scenarios, which, for example, include multiple vehicle classes and driver break times.
Table 8 shows the optimal cost and the average time taken by HGS-CVRP [109, 241] to find this solution, measured over 10 runs with different seeds and a time limit of 10 seconds. Notably, the optimal solution could not be found within this limit, for instance, 21. However, given that HGS-CVRP takes only a few milliseconds to find an optimal solution for typical instances with $n=20$, we expect it to be found within a marginally longer time limit. The optimal solutions were certified using VRPSolver [75, 198, 246, 247].
@@ 4.10.1 Background
Notably, the name of the problem is given by the two fixed parameters. The ODiP has received little attention to date. The DDP has been considered extensively, with the foundations discussed already in the literature from the 1960s [117, 43], and further explored in the 1980s, e.g., [47, 48]. For a comprehensive survey, we refer the reader to [185]. In this work, we consider the ODP, for which–to our knowledge–an NP-hardness proof is not available in the open literature. The ODP arises in the search for network topologies with low latency [225], and a variety of methodologies for constructing networks with low latency are discussed in the literature [24, 217]. We use the ODP variant employed in the GraphGolf competitions [133]: the goal is to solve the ODP, using the average shortest path length (ASPL) as a tie-breaker for graphs with the same diameter, with lower ASPL being preferred. This problem is difficult to solve in practice [133].
@@ 4.10.2 Problem statement
Let $\Delta(s,t)$, $s,t\in V$ be the shortest distance of two vertices in a simple unweighted graph $G=(V,E)$. Given a number $n\in\mathbb{N}$ of vertices, and a maximum vertex degree $d\in\mathbb{N}$. We are looking for a simple graph $G=(V,E)$, $V=\{1,\ldots,n\}$, $E\subseteq\{(u,v)\in V\times V|u\neq v\}$ with $\delta(v)\leq d$, for all $v\in V$, that minimizes $\max_{s,t\in V}\Delta(s,t)$. Multiple formulations for the ODP are possible; for an overview, we refer to [248]. There is considerable freedom in the choice of how the lengths of the all-pairs shortest paths are computed.
@@ 4.10.3 Instances
The instances range from $15$ vertices and degree $3$ to $10^{6}$ vertices and degree $32$. A detailed presentation of the provided instances can be found in Table 9. We provided a program to check solutions, i.e., given a triple $(n,d,k)$ and a graph, it verifies that the graph has $n$ nodes, a maximum degree of $d$, and a diameter of at most $k$.
Figure 16: (Topology) Absolute MIP gap corresponding to order and degree of the flow (top) / quadratic (bottom left) and linearized (bottom right) Seidel model. The size of the points indicates the running time–the bigger, the longer. Furthermore, the timeout is set to 7200s.
@@ 4.10.4 Classical Baseline
Our baseline was computed using Gurobi 11.0.0 on an AMD EPYC-7542 32-core processor using 64 threads with a time limit of 7200s. In the repository, we provide three different MIP formulations. The first model uses a standard s-t flow model to compute shortest path values. The second model uses Seidel’s quadratic all-pairs shortest path formulation. The third model is a linearization of Seidel’s quadratic model. This problem is hard to model as a compact integer program. Gaps and runtimes can be found in Figure 16. The results indicate that we can prove optimality only on rather small instances if we use an MIP solver. With the provided models, even solving $n=30,d=6$ takes too long. However, as can be seen on [133], using heuristic methods, it is possible to compute solutions to larger instances.
@@ 5 Submission Guidelines and Illustrative Results
The following section presents guidelines for submitting benchmark results to [202] and provides exemplary baseline results for three selected problem classes–Low Autocorrelation Binary Sequences, Minimum Birkhoff Decomposition and Independent Set problem–each solved with another quantum optimization algorithm. These results can also be found in the [202] repository.
We provide a submission template in Table 10 which should be submitted together with the discovered solution. The metrics were chosen to enable a fair and systematic comparison of solutions and algorithms. More specifically, we ask for the identification of the concrete problem instance and the submitter(s). Given that not every detail can be given in a compact form, submissions ideally refer to a paper or a code repository with further information, such as hyperparameters, additional hardware specifications, software versions, etc. Furthermore, the best objective value found (for optimization problems) as well as the corresponding solution (in addition to the template) should be given. If supported by an algorithm, an a posteriori bound on the optimal objective value can be provided. Next, the chosen modeling approach should be described (e.g., QUBO, ILP, etc.) as well as the resulting number and type of decision variables and (non-zero) coefficients needed to represent the considered problem instance. Furthermore, the submission should provide a brief summary of the complete optimization workflow to facilitate reproducibility. This includes pre-processing, pre-solver, main optimization algorithm, and post-processing, as well as an indication if the algorithm is deterministic or stochastic. Stochastic algorithms are generally recommended to be repeated multiple times. In this case, the number of successful runs that result in feasible solutions or return solutions close to the best found solutions (cf. Table 10 for more details) should be reported. Lastly, the overall runtime and the runtime spent on the various hardware platforms should be provided. In case of multiple repetitions, the average runtime over all repetitions may be reported. Additional information, such as the distribution of the runtimes or potential correlations with solution quality, is encouraged to be described in a corresponding publication or reference. As discussed in Section 3, the runtimes should not include potential queuing times for hardware access. For most problems, we provide solution checkers in [202]. These should be used to verify proposed solutions before a benchmark result is submitted to [202].
@@
Coefficients Type Type of coefficients, such as integer, binary, continuous.
@@
Workflow Description of the optimization workflow: pre-processing, pre-solvers, machine learning training, optimization algorithms, post-processing, etc.
@@
Hardware Specifications Specifications of hardware used to run the workflow.
Total Runtime (Wall-clock) Total runtime to run the complete workflow.
@@
Table 10: The table summarizes the metrics that should be reported when submitting benchmark results to [202]. All runtimes should be reported as average if multiple algorithm runs were executed.